Calculate the percentage of Oxygen in Vitamin C


1
Expert's answer
2022-04-11T15:38:03-0400

Vitamin C is the ascorbic acid with chemical formula: C6H8O6;

M(C6H8O6) = 176 g/mol;

X(O) = n * Ar(O)/M(C6H8O6) = 6 * 16/176 = 96/176 = 0.5455 = 54.55 %.

Answer: 54.55 %.